# NEAT: Neighborhood-Guided, Efficient, Autoregressive Set Transformer for 3D Molecular Generation

### v1

@@ A.3 Additional Results & Comparisons
In Table 3, we compare NEAT to six diffusion models (EDM (Hoogeboom2022EDM), MiDi (Vignac2023MiDi), GeoLDM (Xu2023GeoLDM), GeoBFN (Song2024GeoBFN), SymDiff (Zhang2025SymDiff), UniGEM (Feng2025UniGEM)) in addition to the autoregressive baselines.
Table 3: Performance comparison of NEAT with diffusion (middle block) and autoregressive (lower block) models trained on the QM9 dataset. All values are given in percent. The best and second-best performing models are highlighted. Models evaluated using our pipeline are marked with a star (*). Metrics are reported as the mean and 95% confidence interval over three runs, each generating 10,000 molecular designs. Results in the top block are computed on the QM9 test set of 13,083 molecules.
metrics
atom
stab.
mol.
stab.
lookup
valid
lookup
unique
xyz2mol
valid
xyz2mol
unique
QM9 99.4 95.3 97.7 97.7 100 100
EDM 98.7 82.0 91.9 90.7 86.7 86.0
MiDi 97.9 84.0 97.9 97.0 n.a. n.a.
GeoLDM 98.9 89.4 93.8 92.7 91.3 90.3
GeoBFN 99.3 93.3 96.9 92.4 n.a. n.a.
SymDiff 98.9 89.7 96.4 94.1 92.8 91.4
UniGEM 98.9 93.2 95.0 93.2 n.a. n.a.
G-SchNet 95.7 68.1 85.5 80.3 75.0 72.5
Symphony 90.8 43.9 68.1 66.5 83.5 81.8
Geo2Seq 98.9 93.2 97.1 81.7 n.a. n.a.
QUETZAL* 98.3 87.5 94.6 90.1 99.3 94.4
$\pm$ 0.1 $\pm$ 0.1 $\pm$ 0.1 $\pm$ 0.2 $\pm$ 0.1 $\pm$ 0.1
NEAT* 98.2 84.7 93.4 88.6 98.5 94.8
$\pm$ 0.1 $\pm$ 0.4 $\pm$ 0.3 $\pm$ 0.4 $\pm$ 0.1 $\pm$ 0.1
Generated Molecules
Figure 5 and 6 show 3D and 2D plots of 35 randomly selected molecules generated by NEAT after being trained on the QM9 dataset.
Figure 5: Randomly selected molecules generated by NEAT after being trained on the QM9 dataset (white: hydrogen, gray: carbon, blue: nitrogen, red: oxygen, green: fluorine). 2D plots of the same molecules are shown in the next figure.
@@
Figure 6: Randomly selected molecules generated by NEAT after being trained on the QM9 dataset (white: hydrogen, gray: carbon, blue: nitrogen, red: oxygen, green: fluorine). Implicit hydrogen atoms have been removed for the sake of clarity. 3D plots of the same molecules are shown in the previous figure.

### v2

@@ A.5 Additional results
We compare NEAT to six diffusion models (EDM (Hoogeboom et al., 2022), GeoLDM (Xu et al., 2023), GCDM (Morehead and Cheng, 2024), GeoBFN (Song et al., 2024), SymDiff (Zhang et al., 2025), UniGEM (Feng et al., 2025)) in addition to the five autoregressive baselines already discussed in the main body of this work. Regarding QM9, we provide both EDM lookup-table and xyz2mol metrics. For GEOM-Drugs, existing literature primarily reports EDM lookup-table metrics, and uniqueness and novelty are often omitted. Due to the unavailability of pre-trained weights for several baseline diffusion models, we report performance values directly from their original publications rather than performing independent re-evaluations. Consequently, due to this lack of available data, some metrics are missing in our comparison in Table 6. Given the observations discussed in Section A.4, both metrics should be interpreted cautiously. Overall, diffusion models slightly outperform autoregressive models on these benchmarks. Nevertheless, as detailed in the main paper, autoregressive models offer practical advantages like adaptive atom counts, faster inference, and straightforward prefix conditioning.
Table 6: Performance comparison of NEAT with diffusion (middle block) and autoregressive (lower block) models trained on the QM9 and GEOM dataset, respectively. All metrics are reported as percentages. The first row reports results on the QM9 and GEOM test set. All other rows report results computed from 10,000 generated molecules. The best and second-best entries are shown in bold and underlined, respectively. Entries for which generated molecules could be retrieved to be re-evaluated with our pipeline are marked with a single star (*). Entries for which the source code and model weights were available to generate molecules are marked with two stars (**). For these models and for NEAT, we report the mean and 95% confidence interval over three runs. Entries marked by a dagger ($\dagger$) were taken from Cheng et al. (2025). Metrics for all other entries were taken from the original publications. xyz2mol entries marked with a double dagger ($\ddagger$) are sourced from publications where the RDKit version is either v2023.3.3 or unspecified; these scores may decrease when re-evaluated with RDKit v2025.9.1.
QM9 GEOM
atom stab. mol. stab. lookup valid lookup unique xyz2mol valid xyz2mol unique atom stab. lookup valid
Test data 99.3 94.8 97.5 97.5 100 100 86.2 96.6
EDM 98.7 82.0 91.9 90.7 86.7†,‡ 86.0†,‡ 81.3 92.6
GeoLDM 98.9 89.4 93.8 92.7 91.3†,‡ 90.3†,‡ 84.4 99.3
GCDM 98.7 86.0 94.9 93.4 n.a. n.a. 88.1 95.5
GeoBFN 99.3 93.3 96.9 92.4 n.a. n.a. 86.2 91.7
SymDiff 98.9 89.7 96.4 94.1 92.8†,‡ 91.4†,‡ 86.2 99.3
UniGEM 99.0 89.8 95.0 93.2 n.a. n.a. 85.1 98.4
G-SchNet* 93.7 63.6 80.1 74.7 73.0 70.6 n.a. n.a.
G-SphereNet* 67.8 14.0 16.4 3.8 37.7 7.2 n.a. n.a.
Symphony* 90.8 43.9 68.1 66.5 77.0 75.2 n.a. n.a.
Geo2Seq 98.9 93.2 97.1 81.7 n.a. n.a. 82.5 96.1
QUETZAL** 98.3 87.5 94.6 90.1 93.7 89.1 86.7 95.7
$\pm$ 0.1 $\pm$ 0.1 $\pm$ 0.1 $\pm$ 0.2 $\pm$ 0.3 $\pm$ 0.2 $\pm$ 0.1 $\pm$ 0.1
NEAT (ours) 98.1 84.2 93.9 88.8 94.4 90.5 81.1 97.9
$\pm$ 0.1 $\pm$ 0.3 $\pm$ 0.1 $\pm$ 0.1 $\pm$ 0.1 $\pm$ 0.2 $\pm$ 0.1 $\pm$ 0.2
Prefixes
The complete list of prefixes extracted from the GEOM training set is shown in Figure 12, 13, and 14.
Vector field visualization
Figures 1517 illustrate iterative molecular construction by the NEAT model trained on QM9. Each row corresponds to one construction step. From left to right, panels show Gaussian-initialized points (orange) advected by the NEAT velocity field (gray arrows, with length indicating speed). Trajectories are integrated using explicit Euler with a uniform time step for $N=60$ steps. The green diamond marks the point selected as the atom position for the next iteration. We use QM9 for visualization, as examples from the larger GEOM dataset become visually cluttered.
@@
Figure 8: Randomly selected molecules generated by NEAT trained on the QM9 dataset. Implicit hydrogen atoms have been removed for the sake of clarity. 3D plots of the same molecules are shown in the next figure.
Figure 9: Randomly selected molecules generated by NEAT trained on the QM9 dataset (white: hydrogen, gray: carbon, blue: nitrogen, red: oxygen, green: fluorine). 2D plots of the same molecules are shown in the previous figure.
Figure 10: Randomly selected molecules generated by NEAT trained on the GEOM-Drugs dataset. Implicit hydrogen atoms have been removed for the sake of clarity. 3D plots of the same molecules are shown in the next figure.
Figure 11: Randomly selected molecules generated by NEAT trained on the GEOM-Drugs dataset (white: hydrogen, gray: carbon, blue: nitrogen, red: oxygen, yellow: sulfur, cyan: fluorine, green: chlorine, brown: bromine). 2D plots of the same molecules are shown in the previous figure.
Figure 12: Prefixes used for the prefix completion task (1/3).
Figure 13: Prefixes used for the prefix completion task (2/3).
Figure 14: Prefixes used for the prefix completion task (3/3).
Figure 15: Iterative build of a molecule with NEAT, trained on QM9 (1/3).
Figure 16: Iterative build of a molecule with NEAT, trained on QM9 (2/3).
Figure 17: Iterative build of a molecule with NEAT, trained on QM9 (3/3).